# LLM2CLIP: Powerful Language Model Unlocks Richer Visual Representation

### v3

@@ 4.2.2 Cross-Lingual Retrieval
Evaluation Dataset.
For Chinese retrieval tasks, we used FlickrCN [18] and CNCOCO [21], which are translated versions of Flickr30K and MS COCO-1K, respectively. LLAVA training and evaluation were conducted using the standard splits from LLAVA 1.5 [24].
LLM Enables English CLIP to Learn Chinese Without Exposure to Chinese Data.
To test LLM knowledge transfer, we conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. As shown in Table 3, EVA02 CLIP and JinaCLIP showed near-zero performance, while LLM2CLIP achieved remarkable results, even outperforming models trained on extensive Chinese data like Wukong [12] and CN-CLIP [42]. This demonstrates the power of LLMs in transferring knowledge to cross-modal spaces.
@@ 4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
Evaluation Dataset.
We conduct zero-shot retrieval testing on Imagenet, following the CLIP method. We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
Imagenet Performance Better In Linear Probe.
Our method shows a slight drop in performance on Imagenet for zero-shot classification. We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions. However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement. This demonstrates that LLM2CLIP can indeed provide better visual features. One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
@@ 4.2.4 Multimodal Large Language Models Performance
Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
LLM2CLIP Improves VLLM Performance.
@@
As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
4.3 Main Conclusions
Directly Replacing CLIP’s Text Encoder with a Vanilla LLM is Harmful.
Directly replacing the text encoder of EVA02 ViT-L/14 with Llama3-8B significantly degrades retrieval performance, as shown by the red-highlighted numbers in Table 7 and Table 6, with average performance drops of 10%-20%. Additionally, the CRA score of the vanilla Llama3-8B is particularly low, indicating that its output features lack sufficient discriminability for captions. These results clearly demonstrate that such an approach imposes a significant burden on CLIP’s learning process.
Improving the Discriminability of LLM Output Features is Key to Integrating LLMs with CLIP.
To enhance the discriminability of LLM output features, we applied Caption Contrastive fine-tuning. Llama3-8B-TC and Llama3-8B-CC represent the results of using pure text corpora(same as LLM2Vec) and a mix of text and augmented CC3M corpora from ShareCaptioner, respectively, both trained with supervised SimCSE loss. As shown in the Table 6, the contrastive learning on mixed caption corpora yields higher CRA scores compared to training on generic text, with noticeable improvements across almost all benchmarks. This difference stems from the fact that LLMs trained on caption-distributed data have stronger discriminability for caption features. These findings highlight the importance of text model discriminability for CLIP training and underscore the rationale behind the caption contrastive fine-tuning in LLM2CLIP.
LLM’s Built-in Knowledge Directly Benefits CLIP.
In Table 7, we observe that directly using LLM2CLIP leads to significant performance improvements even with short caption datasets, which is quite challenging, as CLIP has already seen most of this training data. However, the results from LLM2CLIP with +Llama3-8B-CC still show a clear performance gain, which can be attributed purely to the improvement in the text encoder. On the other hand, +Jina-Bert fails to achieve similar results, and in fact, leads to a performance decline.
LLM Improves CLIP’s Ability to Handle Long and Dense Captions.
Thanks to the LLM’s inherent long context windows and strong text understanding capabilities, CLIP can process long caption datasets more effectively without requiring any architectural modifications. Jina CLIP, for example, enables long-text input by replacing its text encoder with Jina BERT. To verify the LLM’s natural advantage with long texts, we compared our method to Jina CLIP to assess whether the LLM provides a better understanding of long captions. As shown in Table 7, when dense captions are added during training, LLM2CLIP demonstrates significantly greater performance improvements than Jina CLIP, proving that LLMs are better suited to processing long and dense captions.
Larger LLMs Lead to Better Performance.
LLM2CLIP explores the method of replacing CLIP’s text encoder with an LLM and continuing training. We thoroughly investigate the performance differences that various versions of LLMs bring to CLIP. In the experiments shown in Table 6, we swapped the text encoder in CLIP with different language models to explore their impact. We maintained the original LLM2CLIP design, with the only change being the initialization of the text encoder. As expected, within the same Llama family, the 8B model significantly outperfoWrms the 1B model. Mistral Memo also shows a slight improvement over the 8B model. It’s important to note that during CLIP training, the LLM’s gradients remain frozen, further proving that larger LLMs, with their richer knowledge, can deliver substantial performance improvements to CLIP.
4.4 Ablation
Impact of Different Data Ratios.
With the integration of an LLM, LLM2CLIP enhances our ability to understand dense captions. In  Table 8, we present an ablation study that examines the impact of different ratios of dense captions versus original captions on CLIP’s training performance. The “Ratio” column represents the proportion of dense captions used during training. Overall, dense captions contribute to a noticeable improvement in performance, but more is not always better. For example, when using 0% dense captions, the performance on datasets like COCO and Flickr30K remains decent, but long-text benchmarks such as ShareGPT4V, Urban-1K, and DOCCI show poor results. Conversely, with 100% dense captions, performance on short-text retrieval benchmarks is the worst. It’s important to note that the dense captions we used were generated by the ShareCaptioner model, and there may be some distribution differences and noise compared to real data captions, which could somehow explain why using 100% dense data is suboptimal. Our findings indicate that the best performance is achieved when dense captions make up 50% to 75% of the training data.
Impact of Different Data Size.
As shown in  Table 2, larger training datasets consistently yield positive results for LLM2CLIP, with clear improvements in both long-text and short-text retrieval tasks. Our 60M dataset version has already pushed the limits of what CLIP could previously achieve. Even the 3M version, though lightweight, still delivers significant performance gains, demonstrating the efficiency of the LLM2CLIP approach. It’s worth noting that training on the 3M dataset takes only about 3 hours on an 8 H100 GPUs machine, yet results in a transformative leap in performance for a well-pretrained CLIP ViT model.
Adaptor design.
In Table 9, we explore the impact of the Adaptor and its layer count. We find that adding at least one layer of the Adaptor is beneficial, resulting in an average performance improvement of two points. However, increasing it to four layers provides minimal additional gains.
Caption Contrastive Fine-tuning Ablation.
In Table 10, we explore the impact of the key steps in LLM2Vec’s Caption Contrastive fine-tuning, namely MNTP and bidirectional attention. Through comparisons, we observe that MNTP offers almost no positive effect, while bidirectional attention provides some performance gains but is not decisive. Therefore, we conclude that the most significant contribution comes from the supervised SimCSE contrastive learning applied to the captions.
5 Limitations and Future Work
This paper focuses solely on the impact of LLMs on CLIP. In reality, there are many other areas where text can play a transformative role, such as diffusion models [34, 4, 39], Segment Anything [16] and GroundingDINO [25], both of which could potentially see significant performance improvements by applying our method.
While LLMs possess open-world knowledge, aligning them with CLIP may require specific data to fully unlock their potential. LLM2CLIP is a method that requires only a small amount of data for fine-tuning in conjunction with a pretrained CLIP visual encoder, and we have already demonstrated significant performance gains. However, we did not intentionally select fine-tuning data based on specific characteristics. By focusing on aspects such as data distribution, length, or categories, we could further tailor the LLM to address CLIP’s limitations, allowing the LLM to act as a more comprehensive teacher for various tasks.
@@ 6 Conclusion
Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation learning, allowing language — the carrier of human knowledge — to play a pivotal role in spreading the intelligence of LLMs across multiple modalities and applications. The core contributions of this work are threefold: First, we identified the primary reason why LLMs struggle to directly participate in multimodal representation learning — the lack of discriminability in their output features. Second, we introduced the caption contrastive fine-tuning method, which significantly improves feature discriminability, removing the biggest obstacle to utilizing LLMs in CLIP training. Third, we designed the LLM2CLIP training framework, which has proven to be an efficient and effective approach for significantly enhancing the performance of pretrained CLIP models.

### v4

@@ 4.2.2 Cross-Lingual Retrieval
LLMs naturally enable strong multilingual capabilities. As in Table 3, We conducted cross-lingual retrieval experiments where CLIP vision encoders were trained exclusively on English text. We successfully endowed EVA02 with multilingual capabilities, which it previously lacked, and significantly enhanced the already robust multilingual performance of SigLip2.
@@ 4.2.3 Zero-shot Imagenet Classification and Linear Probe Performance
Evaluation Dataset. We conduct zero-shot retrieval testing on Imagenet, following the CLIP method. We also perform linear probing, fine-tuning the model with the exact same hyperparameters: batch size of 1024, learning rate of 0.1, momentum 0.9, weight decay 0 and total 50 epochs using SGD.
Imagenet Performance Better In Linear Probe. Our method shows a slight drop in performance on Imagenet for zero-shot classification. We hypothesize that this is due to insufficient alignment of the data, resulting in some categories having very low word frequency, as well as the LLM’s limited ability to distinguish short captions. However, in the linear probe experiment, we observe that after supplementing the category information, LLM2CLIP actually leads to a greater improvement. This demonstrates that LLM2CLIP can indeed provide better visual features. One of the current limitations is that scaling the dataset to a larger size might be necessary to resolve the zero-shot classification issue.
@@ 4.2.4 Multimodal Large Language Models Performance
Evaluation Dataset. Following LLAVA1.5 [24], we used OpenAI’s CLIP-ViT-L-336 encoder with a simple MLP head to connect to Vicuna-7B. Pretraining included 558K image-caption pairs and 665K visual instruction samples. We fine-tuned the LLAVA1.5 encoder using LLM2CLIP to assess improvements in visual feature extraction.
@@
LLM2CLIP Improves VLLM Performance. As shown in Table 5, LLM2CLIP fine-tuning enhanced multimodal model performance on over 87.5% of benchmarks, with minor losses in only two tasks. This highlights the potential of LLM2CLIP for improving complex image reasoning and understanding tasks.
4.3 Ablation Study: Stage-1 Caption Contrastive Fine-tuning
Experimental Settings. For simplicity, we select Llama 3.1 1B as the default LLM to investigate different design choices for caption contrastive fine-tuning in this section. We use a subset of the CC 3M dataset for caption contrastive fine-tuning. Additionally, we explore augmenting the training data with the Wikitext-103 dataset [25] for MNTP training and the E5 dataset [29] as supplementary pure-text data for caption contrastive fine-tuning.
Structural Design Ablation. We examine various adaptor structures in Table 8. Comparing rows 4 and 5, as well as rows 6 and 7, we observe that neither Linear nor Transformer adaptors significantly impact the performance of subsequent CLIP post-training. Therefore, we default to using no adaptor structure in Stage 1.
Training Method Ablation. We conduct an ablation study of Stage-1 LLM caption contrastive fine-tuning methods in Table Table 9, observing the following key insights: 1. Freezing LLM gradients and relying solely on adaptors yields poor performance, indicating that enabling LoRA is essential. 2. SimCSE is identified as the most critical loss function. MNTP does not significantly enhance SimCSE, and relying solely on MNTP without SimCSE severely deteriorates performance. 3. Supervised SimCSE substantially outperforms unsupervised SimCSE. 4. The performance difference between causal and bidirectional attention mechanisms is negligible.
Ablation for different LLMs. In Table 6, our method achieves significant improvements over these encoders. The result highlighted in red shows that using the original Llama3.1-8B without CC fine-tuning yields notably poor performance, negatively affecting the original CLIP.
4.4 Ablation Study: Stage-2 LLM2CLIP
Adaptor Design. As shown in Table 8, we explore adaptor structures in the Stage-2 LLM2CLIP cross-modal pre-training. Comparing Linear (×$\times 4$) and Transformer (×$\times 1$), their performances are relatively similar, and thus, we select the simpler Linear adaptor structure. Furthermore, in our layer-wise ablation of Linear adaptors, we find that increasing the adaptor size indeed improves performance. Specifically, adaptor performance improves progressively when transitioning from no adaptor to 1, 2 and 4-layer.
Training Method Ablation. We examine several methods of integrating LLMs into CLIP post-training in Table Table 10, with implementation details described in Appendix A. By comparing experiments (a)-(b) and (b)-(c), we find that neither jointly training two losses nor aligning text embeddings significantly improves performance. However, experiment (d), which applies contrastive learning on the concatenated encoder outputs, achieves notable improvements. For simplicity, we do not adopt this design by default but suggest it as a promising direction for future work.
Optimization Strategy Ablation and Efficiency Analysis of LLM2CLIP. As shown in  Table 11, we conducted experiments comparing methods such as LoRA training, frozen LLM online training, and offloading on fixed hardware. Experimental results demonstrate that, with identical computational resources, our proposed offloading method can increase the batch size by tenfold and extend training time by fourfold, ultimately achieving higher performance.
Data Ratio Ablation. In Table Table 7, we experiment with various dataset compositions. The results show a clear trend: higher proportions of dense captions enhance performance in long-text image retrieval tasks, whereas lower dense caption ratios improve short-text retrieval performance. We select a 1:1 ratio by default, as it provides the most balanced improvement for both short and long texts.
Interestingly, even training exclusively on real short captions (100%), our model still significantly boosts performance for both short and long-text retrieval tasks. This improvement can be attributed to the incorporation of LLM, given that real captions share the same distribution as the CLIP pre-training data, yet LLM integration still yields substantial gains.
@@ 5 Conclusion
Given the powerful text understanding capabilities of LLMs, we aim to harness their potential to enhance cross-modal representation, allowing language—the carrier of human knowledge—to play a pivotal role in disseminating the intelligence of LLMs across multiple modalities. This paper successfully proposes an efficient method to seamlessly integrate LLMs into CLIP post-training, demonstrating substantial performance improvements on state-of-the-art multimodal models, including CLIP, EVA02, and SigLip2.
Appendix A Implementation details
Details for Caption Contrastive Fine-tuning
We default to using LoRA for training the LLM, with a rank of 16 and $\alpha=32$. Training is conducted on four nodes, each with 8 NVIDIA A100 40GB GPUs, with an effective batch size of 2048 for one epoch, using a maximum sequence length of 512 tokens for 68 hours. The Adam optimizer is used with a learning rate of $2.5\times 10^{-4}$ and a linear learning rate warm-up for the first 300 steps. We employ bfloat16 quantization, gradient checkpointing, and FlashAttention 2 to optimize GPU memory consumption.
Details for LLM2CLIP Cross-modal Post-training
We employ the AdamW optimizer with a total batch size of 4096 and set the learning rate to $1\times 10^{-5}$ with a cosine decay rate of 0.05.
Details of multi encoder.
When using multiple encoders, we need to align CLIP-V, CLIP-T, and LLM embeddings. We design a one-layer linear projector for each encoder, maintaining the dimensionality at 1024. For the concatenation head scenario, we design a 2048-dimensional head placed after the concatenation of (LLM, CLIP-T) embeddings and another after CLIP-V.
Appendix B Additional results
Training Data Ablation.
In Table Table 12, we explore the impact of different data sources and volumes on Stage-1 LLM caption contrastive fine-tuning. Our findings include: 1. Comparing the combination of 3M Set1 and 1.5M E5 with 4.5M Set1 alone, adding supervised pure-text data from E5 does not provide practical benefits to CLIP’s cross-modal post-training performance and may even slightly degrade performance. 2. For datasets sampled differently from the same distribution (Set1 and Set2), the best results are achieved when Stage 1 and Stage 2 utilize distinct datasets. The overall retrieval performance may depend primarily on the total amount of data seen across both stages; for instance, the performance using Set2 in Stage 1 closely matches that of combining Set1 and Set2.
XM3600 Multilingual Results.
Figure 4 plots radar charts for zero‑shot T2I and I2T retrieval across 36 languages on the XM3600 dataset. Although Siglip2 is already pre‑trained on a large, multilingual corpus, augmenting it with LLM2CLIP—leveraging a large language model to enhance linguistic understanding—still yields a substantial boost in cross‑lingual retrieval performance, even when the fine‑tuning is performed exclusively on English data.